OpenStudy (anonymous):

(a+8)/(a-1) +(a+4)/(a+1) - (8a+2)/(a^2-1)

OpenStudy (anonymous):

to solve this , firstly you have to take LCM

OpenStudy (anonymous):

your LCM will be a^2-1.

hero (hero):

Do you want to end up with a huge expression or a small expression?

OpenStudy (anonymous):

I think small.

hero (hero):

If you want to end up with a small expression you'll have to take a few steps that are not the norm.

hero (hero):

You have \[\frac{a + 8}{a-1} + \frac{a+4}{a+1} - \frac{8a + 2}{a^2 - 1}\]

hero (hero):

And you want to reduce the expression as simply as possible without creating overly large numbers that are annoying to deal with or multiplications that will result in quadratic expressions.

hero (hero):

There's a way to avoid such things. What you do is re-write the expressions in the following manner: \[\frac{a - 1 + 9}{a - 1} + \frac{a + 1 + 3}{a + 1} - \frac{8a + 2}{a^2 - 1}\]

hero (hero):

Then split the fractions like so: \[\frac{a - 1}{a - 1} + \frac{9}{a - 1} + \frac{a + 1}{a + 1} + \frac{3}{a + 1} - \frac{8a + 2}{a^2 - 1}\]

hero (hero):

Notice that you can simplify the above so that you have: \[1 + \frac{9}{a - 1} + 1 + \frac{3}{a + 1} - \frac{8a + 2}{a^2 - 1}\] and then: \[2 + \frac{9}{a - 1} + \frac{3}{a + 1} - \frac{8a + 2}{a^2 - 1}\]

OpenStudy (anonymous):

I think i should do something like this : \[\frac{ a+8(a+1) }{ (a+1)(a-1) } ....\]

hero (hero):

Yes you could do that....if you want to multiply and get quadratic expressions.

OpenStudy (anonymous):

O.K.

hero (hero):

But the approach I am taking avoids all that. Watch what happens when I simplify these expressions

hero (hero):

I am also going to multiply top and bottom by the appropriate expressions:

hero (hero):

\[2 + \frac{9(a + 1)}{(a+1)(a - 1)} + \frac{3(a-1)}{(a-1)(a + 1)} - \frac{8a + 2}{a^2 - 1}\] Which simplifies to: \[2 + \frac{9a + 9}{a^2 - 1} + \frac{3a - 3}{a^2 - 1} - \frac{8a + 2}{a^2 - 1}\]

hero (hero):

And will also combine the fractions to get: \[2 + \frac{9a + 9 + 3a - 3 - (8a + 2)}{a^2 - 1} \] \[2 + \frac{9a + 9 + 3a - 3 - 8a - 2}{a^2 - 1}\] \[2 + \frac{9a + 3a - 8a + 9 - 3 - 2}{a^2 - 1}\] \[2 + \frac{(9 + 3 - 8)a + 4}{a^2 - 1}\] \[2 + \frac{4a + 4}{a^2 - 1}\] \[2 + \frac{4(a + 1)}{(a+1)(a-1)}\]

hero (hero):

Now (a+1)/(a+1) cancels: \[2 + \frac{4\cancel{(a + 1)}}{\cancel{(a+1)}(a-1)}\] Leaving just \[2 + \frac{4}{a-1}\]

hero (hero):

We could leave it in that form, as it is the simplest form, however, your instructor might prefer if you express it as one fraction.

hero (hero):

So I will multiply 2 by (a - 1)/(a - 1): \[\frac{2(a -1)}{a-1} + \frac{4}{a-1}\] This yields: \[\frac{2a - 2}{a - 1} + \frac{4}{a-1}\] Which when combined yields: \[\frac{2a + 4 - 2}{a-1}\] \[\frac{2a + 2}{a-1}\] And finally \[\frac{2(a + 1)}{a-1}\]

OpenStudy (anonymous):

Thank you.